OpenStudy (anonymous):

A rock is thrown vertically upward with an initial speed of 100 ft/s. At the same instant another rock is thrown vertically downward from the top of a 280 ft cliff with an initial speed of 40 ft/s. Neglecting air friction,

OpenStudy (anonymous):

Lol....wat we have to find here....is it the time when they both Will meet ryt ?

OpenStudy (anonymous):

yea the time and position!

OpenStudy (anonymous):

so how then we can find the time?

OpenStudy (anonymous):

V = Vu - ( -Vd) = 100 + 40 = 140ft/s

OpenStudy (anonymous):

V = d/t t = d/t t = 280 / 140 = 2s